Maximize z = 2x1 + 3x2
subject to: 2x1 + x2 ≤ 10
9x1 + 3x2 ≤ 15
x1 ≥ 0, x2 ≥ 0

Answer by ikleyn(52908)

.

1)  In the quadrant QI, draw the feasible domain.


    In this problem, it is a triangle adjacent to x- and y-axes and bounded by the line 

        9x1 + 3x2 = 15.     (1)


    Find coordinates of its vertices.

    One vertex is (0,0), the origin of the coordinate system.

    Two other vertices are x- and y-intercepts of the line (2).


    


    Lines  2x1 + x2 = 10 (read) and  9x1 + 3x2 = 15 (green)



2)  Calculate values of the objective function at these three vertices.



3)  Of the three values of the objective function at vertices of the triangle choose maximum value.

    This value is the maximum value of the objective function in the feasible domain.

    The vertex which provides this maximum value, will give you the solution point to the problem.

That's it.

============

To see other similar problems solved by the Linear Programming method,  look into the lesson
    - Solving minimax problems by the Linear Programming method
in this site.

Also,  look into the solutions of other similar problems in the archive to this forum under the links

https://www.algebra.com/algebra/homework/Inequalities/Inequalities.faq.question.1142000.html
https://www.algebra.com/algebra/homework/Sequences-and-series/Sequences-and-series.faq.question.1137172.html
https://www.algebra.com/algebra/homework/Graphs/Graphs.faq.question.1136382.html
https://www.algebra.com/algebra/homework/Coordinate-system/Coordinate-system.faq.question.1134444.html
https://www.algebra.com/algebra/homework/Graphs/Graphs.faq.question.1131906.html
https://www.algebra.com/algebra/homework/coordinate/word/Linear_Equations_And_Systems_Word_Problems.faq.question.1131043.html
https://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.1129285.html
https://www.algebra.com/algebra/homework/Finance/Finance.faq.question.1128383.html
https://www.algebra.com/algebra/homework/Linear-equations/Linear-equations.faq.question.1123217.html
https://www.algebra.com/algebra/homework/playground/test.faq.question.1112482.html
https://www.algebra.com/algebra/homework/Finance/Finance.faq.question.1102103.html